OpenStudy (anonymous):

Evaluate b2c-1 for b = 8 and c = -4.

OpenStudy (anonymous):

b(2 is power of b) and c( -1 is power of c)

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (midhun.madhu1987):

is this the question??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(8)^2*(-4)=64*(-4)=-256

OpenStudy (midhun.madhu1987):

just plug in the values of b and c..

OpenStudy (anonymous):

OH SOREY... 64/4=16

OpenStudy (anonymous):

-16..I MISSED IT UP.....

OpenStudy (anonymous):

well with my answer choices im not sure how. what do i do with -1

OpenStudy (anonymous):

@keepsecure93

OpenStudy (midhun.madhu1987):